# Sublinear Regret for a Class of Continuous-Time Linear–Quadratic Reinforcement Learning Problems

### v2

@@ 3 A Continuous-Time RL Algorithm
This section presents the steps of designing a continuous-time RL algorithm for solving our LQ problem, including function parameterization, policy evaluation and policy gradient. [^1] As will be explained below, policy evaluation is actually not necessary for the LQ problem considered in this paper. However, we still include it in the discussion and the algorithm for future extension to general problems where policy evaluation is generally needed and indeed a crucial step. We will introduce various techniques such as exploration scheduling and projection for deriving the convergence rate of the policy parameter and the regret bound. We will also describe time discretization for final implementation.
@@ 3.1 Function Parameterization
Inspired by (11) and (12), we parameterize the value function with parameters $\boldsymbol{\theta}\in\mathbb{R}^{d}$:
${J}(t,x;\boldsymbol{\theta})=-\frac{1}{2}{k}_{1}(t;\boldsymbol{\theta})x^{2}+{k}_{3}(t;\boldsymbol{\theta}),$ (13)
@@
Note that (12) suggests that the optimal feedback policy is time-dependent, whose variance depends explicitly on $t$. In our parameterization, the time-dependent variance of the Gaussian policies is replaced by a decaying schedule, called an exploration schedule, of $\phi_{2}$ as a function of the number of iterations, to be presented shortly.
Henceforth we assume that there are positive constants $c_{1},c_{2},c_{3}$ such that $1/c_{2}\leq{k}_{1}(t;\boldsymbol{\theta})\leq{c}_{2}$, $|{k}_{1}^{\prime}(t;\boldsymbol{\theta})|\leq{c}_{1}$ and $|{k}_{3}^{\prime}(t;\boldsymbol{\theta})|\leq{c}_{3}$, for any $0\leq t\leq T$. These assumptions are consistent with the fact that the corresponding functions satisfy the same conditions when the model parameters are known.
@@ 3.2 Policy Evaluation
The general continuous-time PE method developed in (Jia and Zhou, 2022a, ) dictates that one first parameterizes the value function $J(\cdot,\cdot;\pi)$ and the policy $\pi$ by (13) and (14) respectively (with a slight abuse of notation), with the corresponding $p^{\pi}(t)=p(t;\boldsymbol{\phi})$, and then updates $\boldsymbol{\theta}$ in an offline learning setting:
$\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}+\alpha\int_{0}^{T}\frac{\partial J}{\partial\boldsymbol{\theta}}(t,x(t);\boldsymbol{\theta})\left[\mathrm{d}J(t,x(t);\boldsymbol{\theta})-\left(\frac{1}{2}Qx(t)^{2}\mathrm{d}t-\gamma{p}(t;\boldsymbol{\phi})\mathrm{d}t\right)\right],$ (15)
@@
Intriguingly, however, our subsequent theoretical proofs indicate that the convergence and regret results depend only on the bounds (i.e. the constants $c_{1}$, $c_{2}$ and $c_{3}$) of the functions $k_{1}$ and $k_{2}$, not on the specific forms of these functions. This feature is due to the special class of LQ control problems we are tackling. As a result, in our numerical experiments we actually fix a value function (or equivalently $\boldsymbol{\theta}$) throughout without updating it.
@@ 3.3 Policy Iteration
$Z_{1}(s)=\int_{0}^{s}$$\left\{\frac{\partial\log\pi}{\partial\phi_{1}}\left(u(t)\mid x(t);\boldsymbol{\phi}\right)\left[\mathrm{d}J\left(t,x(t);\boldsymbol{\theta}\right)-\frac{1}{2}Qx(t)^{2}\mathrm{d}t\right.\right.$ (17)
@@
As discussed earlier, the other parameter, $\phi_{2}$, controls the level of exploration. In our algorithm, we set a deterministic schedule of this parameter which decreases to 0 as the number of iterations grows. Specifically, we set $\phi_{2,n}=\frac{I_{l}}{b_{n}}$ where $I_{l}$ is the identity matrix of dimension $l$ and $b_{n}\uparrow\infty$ is specified in Theorem 1 below. The order of $b_{n}$ in iteration $n$ is carefully chosen along with those of the other hyperparameters, such as the learning rates, in order to achieve the desired sublinear regret bound of the RL algorithm.
@@ 3.4 Projections
Define $\Pi_{K}(x):=\arg\min_{y\in K}|y-x|^{2}$ to be a general projection mapping a point $x$ onto a given set $K$. Let
$K_{\boldsymbol{\theta},n}=$$\left\{\boldsymbol{\theta}_{n}\in\mathbb{R}^{d}\Big{|}1/c_{2}\leq{k}_{1}(t;\boldsymbol{\theta}_{n})\leq{c}_{2},|{k}_{1}^{\prime}(t;\boldsymbol{\theta}_{n})|\leq{c}_{1},|{k}_{3}^{\prime}(t;\boldsymbol{\theta}_{n})|\leq{c}_{3}\right\},$ (18)
$K_{1,n}=$$\left\{\phi_{1,n}\in\mathbb{R}^{l}\Big{|}|\phi_{1,n}|\leq c_{1,n}\right\},$
where $c_{1},c_{2},c_{3}$ are hyperparameters, and $\{c_{1,n}\}$ is an increasing sequence to be specified in Theorem 1 below. Note here the choice of $K_{\boldsymbol{\theta},n}$ is specific to the special class of LQ problems under consideration – it is independent of $n$ as the regret analysis does not rely on the convergence of $\boldsymbol{\theta}_{n}$. For a general problem, $K_{\boldsymbol{\theta},n}$ needs to be an expanding sequence of sets.
With projection the updating rules for $\boldsymbol{\theta}$ and $\phi_{1}$ in (15) and (16) are modified to
$\begin{split}\boldsymbol{\theta}_{n+1}\leftarrow\Pi_{K_{\boldsymbol{\theta},n+1}}\biggl{(}&\boldsymbol{\theta}_{n}+a_{n}\int_{0}^{T}\frac{\partial J}{\partial\boldsymbol{\theta}}(t,x_{n}(t);\boldsymbol{\theta}_{n})\\ &\left[\mathrm{d}J\left(t,x_{n}(t);\boldsymbol{\theta}_{n}\right)-\frac{1}{2}Qx_{n}(t)^{2}\mathrm{d}t+\gamma{p}\left(t,\boldsymbol{\phi}_{n}\right)\mathrm{d}t\right]\biggl{)},\end{split}$ (19)
$\begin{split}\phi_{1,n+1}\leftarrow\Pi_{K_{1,n+1}}\biggl{(}\phi_{1,n}+a_{n}Z_{1,n}(T)\biggl{)},\end{split}$ (20)
@@ 3.5 Discretization
Our approach for continuous-time RL is characterized by carrying out the entire analysis in the continuous-time setting and discretizing time only at the final implementation stage. The iterations in (19) and (20) involve integrals that can be computed only by approximated discretized summations as well as the $\mathrm{d}J$ term that can be approximated by the temporal difference between two consecutive time steps.  [^2] The discretization errors will be taken into consideration in the appendices; see Theorem 3, Theorem 4, and Remark 2. We therefore discretize the interval $[0,T]$ into uniform time intervals of length $\Delta t$, leading to the following schemes:
$\boldsymbol{\theta}_{n+1}\leftarrow\Pi_{K_{\boldsymbol{\theta},n+1}}\biggl{(}$ $\boldsymbol{\theta}_{n}+a_{n}\sum_{k=0}^{\left\lfloor\frac{T}{\Delta t}-1\right\rfloor}\frac{\partial J}{\partial\boldsymbol{\theta}}(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n})\biggl{[}J\left(t_{k+1},x_{n}(t_{k+1});\boldsymbol{\theta}_{n}\right)$ (22)
$-J\left(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n}\right)-\frac{1}{2}Qx_{n}(t_{k})^{2}\Delta t+\gamma{p}\left(t_{k},\boldsymbol{\phi}_{n}\right)\Delta t\biggl{]}\biggl{)},$
$\phi_{1,n+1}\leftarrow$ $\Pi_{K_{1,n+1}}\biggl{(}\phi_{1,n}+a_{n}\sum_{k=0}^{\left\lfloor\frac{T}{\Delta t}-1\right\rfloor}\biggl{\{}\frac{\partial\log\pi}{\partial\phi_{1}}\left(u_{n}(t_{k})\mid t_{k},x_{n}(t_{k});\boldsymbol{\phi}_{n}\right)$ (23)
$\biggl{[}J\left(t_{k+1},x_{n}(t_{k+1});\boldsymbol{\theta}_{n}\right)-J\left(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n}\right)-\frac{1}{2}Qx_{n}(t_{k})^{2}\Delta t$
$+\gamma{p}\left(t_{k},\boldsymbol{\phi}_{n}\right)\Delta t\biggl{]}+\gamma\frac{\partial{p}}{\partial\phi_{1}}\left(t_{k},\boldsymbol{\phi}_{n}\right)\Delta t\biggl{\}}\biggl{)}.$
@@ 3.6 RL-LQ Algorithm
     $\phi_{2,n}$eterministic sequence of $\phi_{2,n}=\frac{I_{l}}{b_{n}}$ specified in Theorem 1.
@@
         Generate action $u_{n}(t_{k})\sim\boldsymbol{\pi}\left(\cdot\mid t_{k},x_{n}(t_{k});\boldsymbol{\phi}_{n}\right)$ following policy (14) .
         Apply action $u_{n}(t_{k})$ and get new state $x_{n}(t_{k+1})$ by dynamic (1).
         Update time $t_{k+1}\leftarrow t_{k}+\Delta t$ and $t\leftarrow t_{k+1}$.
@@
     Update value function parameters $\boldsymbol{\theta}$ using (22).
     Update policy parameter $\phi_{1}$ using (23).
@@ 4 Regret Analysis
This section presents the main result of the paper – a sublinear regret bound of the RL-LQ algorithm, Algorithm 1. For that, we need to first examine the convergence property and convergence rate of the parameter $\phi_{1,n}$, whose analysis forms the theoretical underpinning of the algorithm. Proofs of the results in this section are provided in Appendices A and B.
@@ 4.1 Convergence of $\phi_{1,n}$
Theorem 1.
In Algorithm 1, let the hyperparameters $c_{1}$, ${c}_{2}$, ${c}_{3}$ and $\gamma$ be fixed positive constants. Set
$a_{n}=\frac{\alpha^{\frac{3}{4}}}{(n+\beta)^{\frac{3}{4}}},\;\;b_{n}=1\vee\frac{(n+\beta)^{\frac{1}{4}}}{\alpha^{\frac{1}{4}}},\;\;c_{1,n}=1\vee(\log\log n)^{\frac{1}{6}},$
@@
    for any $n$, $\mathbb{E}[|\phi_{1,n}-\phi_{1}^{*}|^{2}]\leq{c}\frac{(\log n)^{p}(\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}$, for some positive constants ${c}$ and $p$.
These results ensure the convergence of the learned policy. Moreover, it is a prerequisite for deriving the regret bound of Algorithm 1.
4.2 Regret Bound
Denote
$\bar{J}(\phi_{1},\phi_{2})=\mathbb{E}\left[\int_{0}^{T}\left(-\frac{1}{2}Qx^{\pi}(s)^{2}\right)\mathrm{d}s-\frac{1}{2}Hx^{\pi}(T)^{2}\Big{|}x^{\pi}(0)=x_{0}\right],$ (24)
where $\pi=\mathcal{N}(\cdot|\phi_{1}x,\phi_{2})$.
Theorem 2.
Under the assumptions of Theorem 1, applying Algorithm 1 results in a cumulative regret bound over $N$ episodes given by:
$\sum_{n=1}^{N}\mathbb{E}[\bar{J}(\phi_{1}^{*},0)-\bar{J}(\phi_{1,n},\phi_{2,n})]\leq c+cN^{\frac{3}{4}}(\log N)^{\frac{p+1}{2}}(\log\log N)^{\frac{2}{3}},$
where $c>0$ is a constant independent of $N$, and $p$ is the same constant appearing in Theorem 1.
5 Numerical Experiments
5.1 Simulation Setup
In our simulation study, we consider the case where $l=1$ for the control dimension and $m=1$ for the Brownian motion dimension. The controlled system (1) simplifies to the following form:
$\mathrm{d}x^{u}(t)=(Ax^{u}(t)+Bu(t))\mathrm{d}t+(Cx^{u}(t)+Du(t))\mathrm{d}W(t).$ (25)
In addition, we set the model parameters $A,B,C,D,Q,H,x_{0},T$ to be all 1, and set the exploration schedule $b_{n}=0.2(n+1)^{1/4}$. Other sequences such as that of the learning rate $\{a_{n}\}$ are configured according to the assumptions stated in Theorem 1 and Subsection 3.1; for details see Appendix C.2. In each experiment we execute both our proposed Algorithm 1 and the benchmark Algorithm 2 over $N=400,000$ iterations, while we replicate the experiment independently for 120 times to draw statistical conclusions.
5.2 A Modified Model-Based Algorithm
The algorithms proposed in (Basei et al.,, 2022; Szpruch et al.,, 2024) are designed to estimate parameters $A$ and $B$ in the drift term under the constant volatility setting. To compare with our algorithm tailored for state- and control-dependent volatilities, we extend their methods to include estimating the parameters $C$ and $D$. The details of this modified algorithm are described in Appendix C.1.
5.3 Analysis of Numerical Results
Figures 2 and 2 compare the mean-squared convergence rates of $\phi_{1,n}$ for our model-free Algorithm 1 and the model-based Algorithm 2, using a log-log plot of Mean Squared Error (MSE) versus iterations. The fitted linear regression shows our model’s slope of $-0.5$, confirming Theorem 1 and outperforming the model-based benchmark slope of $-0.08$.
Figure 1: Log-log plot of MSE of learned $\phi_{1,n}$ in Algorithm 1
A comparison of regrets between Algorithms 1 and 2 is presented in Figures 4 and 4. The former yields a regret slope of around 0.73, which is close to the theoretical bound stipulated in Theorem 2 and superior to the slope of 0.88 achieved by the latter.
Figure 3: Log-log plot of the regret of Algorithm 1
Figure 4: Log-log plot of the regret of the benchmark algorithm
These experimental results support the theoretical claims and demonstrate the outperformance of our RL-LQ algorithm compared with its model-based counterpart in terms of both the convergence rates of the policy parameters and the regret bounds.
6 Conclusions
There are several limitations in the setting or results of the paper. First, the state is one-dimensional and the quadratic objective functional (2) has no running reward from controls, which are key assumptions needed to simplify our analysis so that it suffices to consider only (time-invariant) stationary policies. While these assumptions are satisfied in some applications (e.g. in portfolio choice), imposing them is far from satisfactory. We hope that the present paper represents the first step towards solving the general LQ problem and some of the ideas here can inspire a more general convergence/regret analysis for continuous-time RL. Second, we are unable to achieve a better sublinear regret, e.g., a square-root one, which is typical in episodic RL algorithms for tabular or linear MDPs. We are not certain whether that is due to our approach or it is more fundamental due to the diffusion nature of the system dynamics. Finally, extending the analysis to non-LQ problems with general function approximations is an enormous open question. All these point to exciting research opportunities in the (hopefully near) future.
Appendix A A Proof of Theorem 1
@@
$\mathrm{d}x_{n}(t)=(Ax_{n}(t)+B^{\top}u_{n}(t))\mathrm{d}t+\sum_{j=1}^{m}(C_{j}x_{n}(t)+D_{j}^{\top}u_{n}(t))\mathrm{d}W_{n}^{(j)}(t),\quad 0\leq t\leq T,$ (26)
@@
Recall $Z_{1}(\cdot)$ defined by (17), and $Z_{1,n}(T)$ defined by (21) as the value of $Z_{1}(T)$ at the $n$-th iteration. The expectation of $Z_{1,n}(T)$ conditioned on the parameters is denoted by

### v4

@@ 3 A Continuous-Time RL Algorithm
This section presents the steps of designing a continuous-time RL algorithm for solving our LQ problem, including function parameterization, policy evaluation and policy gradient. [^4] As will be explained below, policy evaluation is actually not necessary for the specific LQ problem considered in this paper. However, we still include it in the discussion and in the algorithm for future extension to general problems where policy evaluation is generally needed and indeed a crucial step. We will introduce various techniques such as exploration scheduling and projection for deriving the convergence rate of the policy parameter and the regret bound. We will also describe time discretization for final implementation.
@@ 3.1 Function Parameterization
Inspired by (2.11) and (2.12), we parameterize the value function with parameters $\boldsymbol{\theta}\in\mathbb{R}^{d}$:
${J}(t,x;\boldsymbol{\theta})=-\frac{1}{2}{k}_{1}(t;\boldsymbol{\theta})x^{2}+{k}_{3}(t;\boldsymbol{\theta}),$ (3.1)
@@
Note that (2.12) suggests that the optimal feedback policy is time-dependent, whose variance depends explicitly on $t$. In our parameterization, the time-dependent variance of the Gaussian policies is replaced by a decaying schedule, called an exploration schedule, of $\phi_{2}$ as a function of the number of iterations, to be presented shortly. [^5] An alternative approach is to set a temperature decaying schedule, or equivalently take $\phi_{2}$ as an endogenous, learnable parameter based on the expression (2.12). For example, Szpruch et al., (2024) proposes two different temperature scheduling sequences for their model-based algorithms respectively, which determine the variances of the respective policies in the learning processes that eventually impact the regrets of the algorithms. In this paper, we choose to take a shortcut by taking the entire policy variance as a suitable exogenous tuning sequence (as shown in Theorem 4.1), leaving the temperature $\gamma$ as a constant for simplicity.
Henceforth we assume that there are positive constants $c_{1},c_{2},c_{3}$ such that $1/c_{2}\leq{k}_{1}(t;\boldsymbol{\theta})\leq{c}_{2}$, $|{k}_{1}^{\prime}(t;\boldsymbol{\theta})|\leq{c}_{1}$ and $|{k}_{3}^{\prime}(t;\boldsymbol{\theta})|\leq{c}_{3}$, for any $0\leq t\leq T$. These assumptions are consistent with the fact that the corresponding functions satisfy the same conditions when the model parameters are known. For practical implementations, in general we choose ${k}_{1}(\cdot;\boldsymbol{\theta})$ and ${k}_{3}(\cdot;\boldsymbol{\theta})$ to be neural networks that satisfy these conditions.
@@ 3.2 Policy Evaluation
The general continuous-time PE method developed in (Jia and Zhou, 2022a, ) dictates that one first parameterizes the value function $J(\cdot,\cdot;\pi)$ and the policy $\pi$ by (3.1) and (3.2) respectively (with a slight abuse of notation), with the corresponding $p^{\pi}(t)=p(t;\boldsymbol{\phi})$, and then updates $\boldsymbol{\theta}$ in an offline learning setting:
$\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}+\alpha\int_{0}^{T}\frac{\partial J}{\partial\boldsymbol{\theta}}(t,x(t);\boldsymbol{\theta})\left[\mathrm{d}J(t,x(t);\boldsymbol{\theta})-\left(\frac{1}{2}Qx(t)^{2}\mathrm{d}t-\gamma{p}(t;\boldsymbol{\phi})\mathrm{d}t\right)\right],$ (3.3)
@@
Intriguingly, however, our subsequent theoretical analysis indicates that the convergence and regret results depend only on the bounds (i.e. the constants $c_{1}$, $c_{2}$ and $c_{3}$) of the functions $k_{1}$ and $k_{2}$, not on the specific forms of these functions. This feature is due to the special class of LQ control problems we are tackling. As a result, in our numerical experiments we actually fix a value function (or equivalently $\boldsymbol{\theta}$) throughout without updating it. However, we still include the policy evaluation part in our exposition as it is needed in more general cases.
@@ 3.3 Policy Iteration
$Z_{1}(s)=\int_{0}^{s}$$\left\{\frac{\partial\log\pi}{\partial\phi_{1}}\left(u(t)\mid x(t);\boldsymbol{\phi}\right)\left[\mathrm{d}J\left(t,x(t);\boldsymbol{\theta}\right)-\frac{1}{2}Qx(t)^{2}\mathrm{d}t\right.\right.$ (3.5)
@@
As discussed earlier, the other parameter, $\phi_{2}$, controls the level of exploration. In our algorithm, we set a deterministic schedule of this parameter which decreases to 0 as the number of iterations grows. Specifically, we set $\phi_{2,n}=\frac{I_{l}}{b_{n}}$ where $I_{l}$ is the identity matrix of dimension $l$ and $b_{n}\uparrow\infty$ is specified in Theorem 4.1 below. The order of $b_{n}$ in iteration $n$ is carefully chosen along with those of the other hyperparameters, such as the learning rates, in order to achieve the desired sublinear regret bound of the RL algorithm.
@@ 3.4 Projections
First, in general ${k}_{1}(\cdot;\boldsymbol{\theta})$ and ${k}_{3}(\cdot;\boldsymbol{\theta})$ are two fixed function approximators (e.g. neural networks), which can be properly chosen so that there exist constants $c_{1},c_{2},c_{3}$ satisfying
$1/c_{2}\leq{k}_{1}(t;\boldsymbol{\theta})\leq{c}_{2},|{k}_{1}^{\prime}(t;\boldsymbol{\theta})|\leq{c}_{1},|{k}_{3}^{\prime}(t;\boldsymbol{\theta})|\leq{c}_{3},\;\;\;\forall(t,\boldsymbol{\theta}).$
Fix a constant $c_{\boldsymbol{\theta}}>0$ and define
$K_{\boldsymbol{\theta}}=\left\{\boldsymbol{\theta}\in\mathbb{R}^{d}\Big{|}|\boldsymbol{\theta}|\leq c_{\boldsymbol{\theta}}\right\}.$ (3.6)
Note that the definition of $K_{\boldsymbol{\theta}}$ is specific to our special class of LQ problems under consideration – it is independent of $n$ because the subsequent regret analysis, as it turns out, does not rely on the convergence of $\boldsymbol{\theta}$. Importantly, the hyperparameters $c_{1},c_{2},c_{3}$ are determined by the specified function approximators ${k}_{1}(\cdot;\boldsymbol{\theta})$ and ${k}_{3}(\cdot;\boldsymbol{\theta})$, without requiring prior knowledge of the model parameters. (Later in our numerical experiments we will simply set $k_{1}\equiv 1$ and $k_{3}\equiv 0$.)
Next, define
$K_{1,n}=\left\{\phi_{1}\in\mathbb{R}^{l}\Big{|}|\phi_{1}|\leq c_{1,n}\right\},$ (3.7)
where $\{c_{1,n}\}$ is an increasing sequence to be specified in Theorem 4.1. Clearly, $K_{1,n}\subset K_{1,n+1}\subset\cdots$ and $\cup_{n}K_{1,n}=\mathbb{R}^{l}$. Finally, for any nonempty, convex and compact set $K$, we define $\Pi_{K}(x):=\arg\min_{y\in K}|y-x|^{2}$ to be the standard projection mapping of a point $x$ onto $K$. In our updating rule below, we mainly control the learned policy $\boldsymbol{\phi}_{n}\in K_{1,n}$ using projection so that it will not grow too fast to an unbounded region.
With projections, the updating rules for $\boldsymbol{\theta}$ and $\phi_{1}$ in (3.3) and (3.4) are modified as follows. Let $x_{n}$ and $u_{n}$ be the state trajectory and the control trajectory at the $n$-th iteration, obtained from the systems dynamic (2.1) with a stochastic policy (3.2), i.e., $u_{n}(t)|x_{n}(t)\sim\mathcal{N}(\phi_{1,n}x_{n}(t),\phi_{2,n})$. We update
$\begin{split}\boldsymbol{\theta}_{n+1}\leftarrow\Pi_{K_{\boldsymbol{\theta}}}\biggl{(}&\boldsymbol{\theta}_{n}+a_{n}\int_{0}^{T}\frac{\partial J}{\partial\boldsymbol{\theta}}(t,x_{n}(t);\boldsymbol{\theta}_{n})\\ &\left[\mathrm{d}J\left(t,x_{n}(t);\boldsymbol{\theta}_{n}\right)-\frac{1}{2}Qx_{n}(t)^{2}\mathrm{d}t+\gamma{p}\left(t,\boldsymbol{\phi}_{n}\right)\mathrm{d}t\right]\biggl{)},\end{split}$ (3.8)
$\begin{split}\phi_{1,n+1}\leftarrow\Pi_{K_{1,n+1}}\biggl{(}\phi_{1,n}+a_{n}Z_{1,n}(T)\biggl{)},\end{split}$ (3.9)
@@ 3.5 Discretization
Our approach for continuous-time RL is characterized by carrying out the entire analysis in the continuous-time setting and discretizing time only at the final implementation stage. There are two reasons for discretization. First, note that in the updating rules (3.8) to (3.10), we implicitly assume a continuous sampling of control $u_{n}(t)|x_{n}(t)\sim\mathcal{N}(\phi_{1,n}x_{n}(t),\phi_{2,n})$, which is not practical for actual inplementation. Second, The iterations in (3.8) and (3.9) involve integrals that can be computed only by approximated discretized summations as well as the $\mathrm{d}J$ term that can be approximated by the temporal difference between two consecutive time steps.666 Therefore, the error due to discretization also has two parts – one by discretely sampling randomized controls and the other by computing the integrals using finite sums. Szpruch et al., (2024) focuses on the first type of error and Basei et al., (2022) on the second one in their analyses respectively. We address both types of discretization errors in the proof of Theorem 4.1, as well as in Theorems B.3 and B.6 in Appendix B of the full online version of this paper, available at https://arxiv.org/pdf/2407.17226 (which hereafter will be referred to as the “online version”). Therefore, at the $n$th iteration we discretize the interval $[0,T]$ into uniform time intervals of length $\Delta t_{n}$, leading to the following schemes:
$\boldsymbol{\theta}_{n+1}\leftarrow\Pi_{K_{\boldsymbol{\theta},n+1}}\biggl{(}$ $\boldsymbol{\theta}_{n}+a_{n}\sum_{k=0}^{\left\lfloor\frac{T}{\Delta t_{n}}-1\right\rfloor}\frac{\partial J}{\partial\boldsymbol{\theta}}(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n})\biggl{[}J\left(t_{k+1},x_{n}(t_{k+1});\boldsymbol{\theta}_{n}\right)$ (3.11)
$-J\left(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n}\right)-\frac{1}{2}Qx_{n}(t_{k})^{2}\Delta t_{n}+\gamma{p}\left(t_{k},\boldsymbol{\phi}_{n}\right)\Delta t_{n}\biggl{]}\biggl{)},$
$\phi_{1,n+1}\leftarrow$ $\Pi_{K_{1,n+1}}\biggl{(}\phi_{1,n}+a_{n}\sum_{k=0}^{\left\lfloor\frac{T}{\Delta t_{n}}-1\right\rfloor}\biggl{\{}\frac{\partial\log\pi}{\partial\phi_{1}}\left(u_{n}(t_{k})\mid t_{k},x_{n}(t_{k});\boldsymbol{\phi}_{n}\right)$ (3.12)
$\biggl{[}J\left(t_{k+1},x_{n}(t_{k+1});\boldsymbol{\theta}_{n}\right)-J\left(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n}\right)-\frac{1}{2}Qx_{n}(t_{k})^{2}\Delta t_{n}$
$+\gamma{p}\left(t_{k},\boldsymbol{\phi}_{n}\right)\Delta t_{n}\biggl{]}+\gamma\frac{\partial{p}}{\partial\phi_{1}}\left(t_{k},\boldsymbol{\phi}_{n}\right)\Delta t_{n}\biggl{\}}\biggl{)}.$
@@ 3.6 RL-LQ Algorithm
    $\phi_{2,n}$eterministic sequence of $\phi_{2,n}=\frac{I_{l}}{b_{n}}$ specified in Theorem 4.1.
@@
         Generate action $u_{n}(t_{k})\sim\boldsymbol{\pi}\left(\cdot\mid t_{k},x_{n}(t_{k});\boldsymbol{\phi}_{n}\right)$ following policy (3.2) .
         Apply action $u_{n}(t_{k})$ and get new state $x_{n}(t_{k+1})$ by dynamic (2.1).
         Update time $t_{k+1}\leftarrow t_{k}+\Delta t_{n}$ and $t\leftarrow t_{k+1}$.
@@
    Update value function parameters $\boldsymbol{\theta}$ using (3.11).
    Update policy parameter $\phi_{1}$ using (3.12).
@@ 4 Regret Analysis
This section presents the main result of the paper – a sublinear regret bound of the RL-LQ algorithm, Algorithm 3.1. For that, we need to first examine the convergence property and convergence rate of the parameter $\phi_{1,n}$, whose analysis forms the theoretical underpinning of the algorithm.
@@ 4.1 Convergence of $\phi_{1,n}$
Theorem 4.1.
In Algorithm 3.1, let the hyperparameters $c_{1}$, ${c}_{2}$, ${c}_{3}$ and $\gamma$ be fixed positive constants. Set
$a_{n}=\frac{\alpha^{\frac{3}{4}}}{(n+\beta)^{\frac{3}{4}}},\;\;b_{n}=1\vee\frac{(n+\beta)^{\frac{1}{4}}}{\alpha^{\frac{1}{4}}},\;\;c_{1,n}=1\vee(\log\log n)^{\frac{1}{6}},\;\;\Delta t_{n}=T(n+1)^{-\frac{5}{8}}$
@@
    for any $n$, $\mathbb{E}[|\phi_{1,n}-\phi_{1}^{*}|^{2}]\leq{c}\frac{(1\vee\log n)^{p}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}$, for some positive constants ${c}$ and $p$.
Proof.
To keep our proof easy to follow, we first present the analysis by ignoring both types of discretization error pointed out in Section 3.5. More precisely, we first assume a continuous sampling of control $u_{n}(t)|x_{n}(t)\sim\mathcal{N}(\phi_{1,n}x_{n}(t),\phi_{2,n})$ with $\Delta t_{n}=0$ and use updating rules (3.8)–(3.10). After we prove the desired conclusion without discretization error, we address the impact of the discretization error and show that it will not change the final conclusion.
@@
$\mathrm{d}x_{n}(t)=(Ax_{n}(t)+B^{\top}u_{n}(t))\mathrm{d}t+\sum_{j=1}^{m}(C_{j}x_{n}(t)+D_{j}^{\top}u_{n}(t))\mathrm{d}W_{n}^{(j)}(t),\quad 0\leq t\leq T,$ (4.1)
@@
Recall $Z_{1}(\cdot)$ defined by (3.5), and $Z_{1,n}(T)$ defined by (3.10) as the value of $Z_{1}(T)$ at the $n$-th iteration. The expectation of $Z_{1,n}(T)$ conditioned on the parameters is denoted by